The nurse has received an order for penicillin v potassium (Penicillin-VK) suspension 0.75 g. The drug is available as 250 mg/5 ml. How many milliliters should be given?
Answer #1

ANSWER.

* Calculation of amount (ml) of medication required.

       Formula = Desired dose/Dose in hand * Quantity.

   Here

     Desired dose (D) = 0.75 g.

Convert gm value to mg for calculation

                  1g = 1000 mg.

So

     Desired dose(D) = 750 mg.

     Dose in hand (H) = 250 mg.

         Quantity(Q) = 5 ml.

As per formula

    Amount(ml) of drug required = D/H * Q

           = 750/250 * 5

   = 3 * 5

   = 15 ml.